OpenStudy (anonymous):

what is the antiderivative of 5x?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Im sorry it5/2 x^2

OpenStudy (anonymous):

You forget the constant.

OpenStudy (radar):

The derivative of your proposed answer would be 5x not 5.

OpenStudy (radar):

Try (5/2)x+C

OpenStudy (anonymous):

Nice catch, radar.

OpenStudy (radar):

I'm wrong it is 5x+C

OpenStudy (anonymous):

\[\int\limits_{}5xdx=5\int\limits_{}xdx=(5/2)x^2+c\]

OpenStudy (anonymous):

lol. I was looking at the five you originally posted as the equation being integrated, nadeem.

OpenStudy (radar):

I'm confused what is the derivative of (5/2)x^2+C Lets see it is 5x Yeah

OpenStudy (anonymous):

the 5 is a constant and an anitderivative is an integral

OpenStudy (anonymous):

yes the derivative of (5/2)x^2+C = 5x... check

OpenStudy (anonymous):

Well take the derivative of your proposed antiderivative. Nadeem had it right the second time.